LIF SPECTROSCOPY OF JET-COOLED CHD_{2}O

\underline{XIAOYONG LIU}, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210; ILIA J. KALINOVSKI^{a}, Chemical Sciences Division of the Lawrence Berkeley National Laboratory and Department of Chemistry of UC Berkeley, California 94720; and C. BRADLEY MOORE, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210 and Chemical Sciences Division of the Lawrence Berkeley National Laboratory and Department of Chemistry of UC Berkeley, California 94720.

Rotationally resolved Laser-induced Fluorescence measurement is performed on the electronic transition Ã^{2}A′ ← X̃^{2}E_{3/2} of jet-cooled CHD_{2}O from 2900-3100 cm^{−1} in vibrational excitation energy of the Ã^{2}A′ state, where ab initio result suggests the fundamental of C-H stretch is located. Resorting to a rotational analysis, eight vibrational bands are recognized and characterized. The vibrational assignments are confirmed by a relative intensity analysis and the symmetries thoughtfully determined in the rotational analysis.

aNovellus Systems, Inc., 2730 Junction Ave., San Jose, CA 95134